Find graphically and analytically the resultant of the following displacements A 10m northwest B 20m30° north of east C 35m due south

1
Expert's answer
2021-12-30T11:53:59-0500

"A_N=10\cos45°=7.1\text{ m}.\\
A_E=-10\sin45°=-7.1\text{ m}.\\
B_N=20\sin30°=10\text{ m}.\\
B_E=20\cos30=17\text{ m}.\\
C_N=-35\text{ m},\\
C_E=0.\\\space\\
R_N=A_N+B_N+C_N=-17.9\text{ m},\\
R_E=A_E+B_E+C_E=9.9\text{ m},\\
R=\sqrt{R_N^2+R_E^2}=20\text{ m}.\\\space\\
\theta=\arctan\dfrac{R_E}{R_N}=28.9°\text{ E of S}."